OpenStudy (anonymous):

Put the following equation in standard form. 6y = 8x +1/4

OpenStudy (anonymous):

Put it in this form: \[Ax+By=C \quad A,B,C \in \mathbb{Z}\]

OpenStudy (anonymous):

8x-6y+1/4=0 or 32x-24+1=0..

OpenStudy (anonymous):

^That's wrong.

OpenStudy (anonymous):

or 8x-6y=-1/4 or 32x+24y= -1

OpenStudy (anonymous):

Standard form of a line is y=mx+c

OpenStudy (anonymous):

@beeqay, I'm honestly not sure which standard form the question refers to. Typically, \(Ax+By=C\) is the standard form. (A, B, and C are integers.)

OpenStudy (anonymous):

Wikipedia lists it here: http://en.wikipedia.org/wiki/Linear_equation#Standard_form

OpenStudy (anonymous):

@Limitless Yes but that is a straight line. While your equation is true, I think that is true for parabolas.

OpenStudy (anonymous):

..Parabolas? Those are quadratic equations. This is a linear equation. What are you talking about?

OpenStudy (anonymous):

What's the shape of a quadratic equation in graph?

OpenStudy (anonymous):

y=mx+b is an slope intercept form..

OpenStudy (anonymous):

Coolstude, I said you were wrong because you didn't do what the question asked. You also posted an answer, rather than helping.

OpenStudy (anonymous):

Beeqay, it depends entirely on what your quadratic is. This is related to the study of conics.

OpenStudy (anonymous):

hmm, actually yea that is true. ax+by=c is the standard form Limitless, a quadratic equation is a parabola, hence my reference to parabolas. We on the same page now?

OpenStudy (anonymous):

Yeah, actually. I wasn't aware that all quadratics were strictly parabolas. However, I'm not entirely sure what parabolas have to do with this.

OpenStudy (anonymous):

Yea never mind, I thought the standard form of a line would be it's slope intercept line, been too too long since I saw such a question.